# Modeling non-genetic information dynamics in cells using reservoir computing

### 13 Dec 2023 v1

@@ Acknowledgement
We thank Dr. Robert Eil (University of Oregon) for providing flow cytometry data for this work. DN acknowledges Dr. Prabesh Khatiwada (Columbia University) for insightful discussion on cellular immune response and flow cytometry.
@@
![[Uncaptioned image]](extracted/5290580/point_source_noise_cos_to_sin_noise_level_0.05.png)
@@
![[Uncaptioned image]](extracted/5290580/point_source_noise_cos_to_sin_noise_level_0.1.png)
@@
![[Uncaptioned image]](extracted/5290580/point_source_noise_cos_to_sin_noise_level_0.25.png)
@@
![[Uncaptioned image]](extracted/5290580/point_source_noise_cos_to_sin_noise_level_0.5.png)
@@
![[Uncaptioned image]](extracted/5290580/point_source_noise_cos_to_sin_noise_level_1.0.png)
@@
![[Uncaptioned image]](extracted/5290580/spherical_source_noise_cos_to_sin_noise_level_0.05.png)
@@
![[Uncaptioned image]](extracted/5290580/spherical_source_noise_cos_to_sin_noise_level_0.1.png)
@@
![[Uncaptioned image]](extracted/5290580/spherical_source_noise_cos_to_sin_noise_level_0.25.png)
@@
![[Uncaptioned image]](extracted/5290580/spherical_source_noise_cos_to_sin_noise_level_0.5.png)
@@
![[Uncaptioned image]](extracted/5290580/spherical_source_noise_cos_to_sin_noise_level_1.0.png)

### 14 Mar 2024 v2

@@ Acknowledgement
We thank Dr. Robert Eil (Oregon Health & Science University) for providing flow cytometry data for this work. DN acknowledges Dr. Prabesh Khatiwada (Columbia University) for insightful discussion on cellular immune response and flow cytometry.
@@
![[Uncaptioned image]](extracted/5471664/point_source_noise_cos_to_sin_noise_level_0.05.png)
@@
![[Uncaptioned image]](extracted/5471664/point_source_noise_cos_to_sin_noise_level_0.1.png)
@@
![[Uncaptioned image]](extracted/5471664/point_source_noise_cos_to_sin_noise_level_0.25.png)
@@
![[Uncaptioned image]](extracted/5471664/point_source_noise_cos_to_sin_noise_level_0.5.png)
@@
![[Uncaptioned image]](extracted/5471664/point_source_noise_cos_to_sin_noise_level_1.0.png)
@@
![[Uncaptioned image]](extracted/5471664/spherical_source_noise_cos_to_sin_noise_level_0.05.png)
@@
![[Uncaptioned image]](extracted/5471664/spherical_source_noise_cos_to_sin_noise_level_0.1.png)
@@
![[Uncaptioned image]](extracted/5471664/spherical_source_noise_cos_to_sin_noise_level_0.25.png)
@@
![[Uncaptioned image]](extracted/5471664/spherical_source_noise_cos_to_sin_noise_level_0.5.png)
@@
![[Uncaptioned image]](extracted/5471664/spherical_source_noise_cos_to_sin_noise_level_1.0.png)